# The Cycle Counts of Graphs

### v2

@@ Abstract.
We prove that an inseparable graph can have any positive number of cycles with the six exceptions 2, 4, 5, 8, 9, 16, and that an inseparable cubic graph has the sole additional exception 13.
@@
The only positive integers that are not cycle count numbers are $2$, $4$, $5$, $8$, $9$, and $16$.
Since the number of cycles is unchanged upon subdividing edges, or the reverse operation, Theorem 1 is equally true for multigraphs and simple graphs. However, by “graph” we mean to allow multiple edges and loops; results do not necessarily hold for simple graphs. The following is an example.
@@
The only positive integers that are not cycle counts of inseparable cubic graphs are $1$, $2$, $4$, $5$, $8$, $9$, $13$, and $16$. The same positive integers are the only ones that are not cycle counts of inseparable planar cubic graphs.
These results are Sequences A385523 and A385524 in the Online Encyclopedia of Integer Sequences [4].
@@
Let $c(G)$ be the number of cycles in $G$. This number depends only on the homeomorphism type of $G$, so we may freely introduce new vertices by subdividing edges, or execute the reverse operation of suppressing divalent vertices. We denote by $\widehat{G}$ any graph that is homeomorphic to $G$ (i.e., they have isomorphic subdivisions). All our graphs are assumed to be inseparable, which means connected and having no cutpoint (such as a vertex that supports a loop and is incident with another edge) but not necessarily simple. Inseparability is the same as 2-connectivity, except that the graph with one vertex and one loop is inseparable without being 2-connected.
An ear in a graph is a path (with distinct endpoints) of positive length whose internal vertices, if any, have degree 2, and whose endpoints have degree greater than 2. When adding an ear, we allow ourselves to subdivide edges to become the endpoints of the ear. It is known that every inseparable graph is a cycle or obtained from a cycle by successively adding ears, and that when we add an ear to an inseparable graph, the result is always inseparable ([6, Theorem 19], [5, Theorem III.11]). The opposite action is deleting an ear, which involves removing the edges and internal vertices of the ear but not its endpoints. We assume this is only done to leave an inseparable remaining graph.
@@ 1. Proof for numbers from 1 to 19
The notation $G/S$ means the contraction of $G$ by $S\subseteq E$. A minor of $G$ is any contraction of a subgraph.
@@
We prove a lemma that tells us how to get the most or fewest cycles by adding an ear. A vertex is divalent if its degree is 2, and multivalent if it has degree $>2$. Note that an inseparable $G$ will have at least two multivalent vertices unless it is a cycle.
@@
To prove (c), suppose that $v$ and $w$ lie on the same ear $P$ of $\widehat{G}$, but $v$ is not at the end. As before, contract $P_{xv}$. Then $\widehat{G}\smallsetminus P^{\prime}$ is separable, $w$ being a cutpoint. Therefore, by Proposition  4(c) the number of cycles is unchanged.
@@
Now we prove Theorem 1 for numbers 4 through 7. To have more than 3 cycles we must add an ear to a theta graph. The four possible graphs are shown in Figure 2. They have 6 cycles, except for $K_{4}$ which has 7 cycles. This also shows that counts 4 and 5 are impossible.
@@
Now we see how to add ears three and four to a $\varTheta_{4}^{\prime}$ in a way that never creates a $\widehat{K}_{4}$ subgraph. (If we did create such a subgraph, the result could be treated as a $\widehat{K}_{4}$ with added ear or ears.)
First we add ear three. The $\varTheta_{4}^{\prime}$ is built on a $\varTheta$, whose constituent paths we call $P,Q,R$, by placing ear two on one of the paths, say $P$. Let $P=P_{ux}P_{xy}P_{yv}$, where $P_{ux}$ and $P_{yv}$ may have length 0. (See Figure 2.) Each further ear must also be on one of $P$ as in Figure 4(a2, b) or $Q$ as in Figure 4(a1), or on ear two, which is equivalent to being on $P$ nested inside ear two as in Figure 4(a2). In Figure 4(a) we get $\binom{4}{2}=10$ cycles because its cycle count is the same as that of $\varTheta_{5}$ by Lemma 5(c). In Figure 4(b) we have a necklace of three cycles (it is the lower left graph in Figure 1); this graph has $1+3+2\cdot 3=11$ cycles. We note that adding an ear on a previously added ear does not give any different graphs from these.
@@
Now we add ear four. The four essentially different ways to do so are shown in Figure 5. Graphs (b) show all ways to have ears two, three, and four on the same constituent path of the theta graph; graphs (a) show the other ways. The four types are represented by (a1–a3) and (b1). By Lemma 5(c) graphs (a1) and (a3) have the same cycle count as $\varTheta_{6}$, that is, $\binom{6}{2}=15$. Graph (b1) is a necklace of four small cycles; it has the 4 small cycles and $2^{4}$ large cycles by choosing one edge in each small cycle, a total of 20 cycles. Graph (a2) has 3 small cycles, $2$ cycles using the left and right theta constituent paths, $2^{2}$ using the center and right constituent paths, and $2^{3}$ using the center and left constituent paths, for a total of 17 cycles (in fact by Lemma 5(c) it has the same cycle count as the bottom left graph in Figure 1). It is easy to verify that no way of adding ears on ears gives any other graphs.
@@ 2. Proof for numbers beyond 89
Now we return to the proof of the main theorem. It was shown in [1] that every integer other than 2, 4, 5, 7, 8, 9, 12, 13, 14, 16, 18, 19, 22, 23, 26, 27, 29, 31, 33, 35, 38, 39, 42, 43, 46, 50, 52, 54, 60, 65, 68, 72, 77 and 89 counts the number of subtrees of a tree. (Cf. [4, Sequence A184164].) By Theorem 7, every positive integer other than those (and 1) is the number of cycles in a 2-connected graph.
@@ 3. Proof for the missing numbers up to 89
Figure 7. Graphs for the 22 exceptional cycle counts from 22 to 89.
@@ 4. Proof of Theorem 2
The construction in Section 2 produces inseparable, planar cubic graphs for cycle counts not listed in [1] ([4, Sequence A184164]), as we noted there. This leaves us with the counts listed in [1] apart from 1, 2, 4, 5, 8, 9, 13, and 16. A computer search using NAUTY [3] found planar cubic graphs for these cases, with up to 14 vertices. An example which requires 14 vertices is shown in Figure 8.
@@ 5. Consequences
Every positive integer is the cycle count of some inseparable Hamiltonian planar graph, except for $2$, $4$, $5$, $8$, $9$, and $16$.
@@ 6. Declarations
Data for inseparable planar graphs of order at most 8 having all cycle counts up to 100 are posted at [2] under the heading “Graphs with given cycle counts”.

### v3

@@ Abstract.
We prove that an inseparable graph can have any positive number of cycles with the six exceptions 2, 4, 5, 8, 9, 16, and that an inseparable cubic graph has the additional exceptions 1 and 13. The exceptions for simple inseparable cubic graphs are unknown.
@@
The only positive integers that are not cycle count numbers are $2$, 44, 55, 88, 99, and $16$.
Since the number of cycles is unchanged upon subdividing edges, or the reverse operation, Theorem 1 is equally true for multigraphs and simple graphs. In this article we write “graph” to allow multiple edges but not loops, and “simple graph” when neither loops nor multiple edges are allowed. We will only consider inseparable graphs, which are those without cutpoints. For loop-free graphs with more than two vertices, inseparability is the same as 2-connectivity. The following is an example of a theorem that is true for graphs but not for simple graphs (see Section 6).
@@
The only positive integers that are not cycle counts of inseparable cubic graphs are $1$, 22, 44, 55, 88, 99, $13$, and $16$. The same positive integers are the only ones that are not cycle counts of inseparable planar cubic graphs.
The numbers in Theorems 1 and 2 are sequences A385523 and A385524 in the Online Encyclopedia of Integer Sequences [6].
@@
Let $c(G)$ be the number of cycles in $G$. This number depends only on the homeomorphism type of $G$, so we may freely introduce new vertices by subdividing edges, or execute the reverse operation of suppressing divalent vertices. We denote by $\widehat{G}$ any graph that is homeomorphic to $G$ (i.e., they have isomorphic subdivisions).
An ear in a graph is a path (with distinct endpoints) of positive length whose internal vertices, if any, have degree 2, and whose endpoints have degree greater than 2. When adding an ear, we allow ourselves to subdivide edges to become the endpoints of the ear.
Our approach will use the following famous theorem of Whitney.
Theorem 3 ([7, Theorem 19]).
Adding an ear to an inseparable graph gives an inseparable graph. Moreover, every inseparable graph with more than two vertices is a cycle or can be obtained from a cycle by successively adding ears.
In particular, as we construct inseparable graphs by adding ears, we do not need to add ears to graphs that are separable.
@@ 1. Proof for numbers from 1 to 19
The notation $G/S$ means the contraction of $G$ by $S\subseteq E$, while $G\smallsetminus P$ means $G$ with the edges and internal vertices of the ear $P$ (but not its endpoints) removed. A minor of $G$ is any contraction of a subgraph. If $v,w$ are two vertices of an ear $P$, then $P_{vw}$ denotes the subpath of $P$ whose endpoints are $v$ and $w$.
@@
We prove a lemma that tells us how to get the most or fewest cycles by adding an ear. A vertex is divalent if its degree is 2, and multivalent if it has degree greater than 2. Note that an inseparable $G$ will have at least two multivalent vertices unless it is a cycle.
@@
To prove (c), suppose that $v$ and $w$ lie on the same ear $P$ of $\widehat{G}$, but $v$ is not at the end. As before, contract $P_{xv}$. Since $\widehat{G}\smallsetminus P_{xv}$ is separable, $w$ being a cutpoint, the number of cycles is unchanged, by Proposition 5(c).
@@
Now we prove Theorem 1 for numbers 4 through 7. To have more than 3 cycles we must add an ear to a theta graph. (Recall from Theorem 3 that we don’t need to consider adding ears to separable graphs.) The four possible graphs are shown in Figure 1. They have 6 cycles, except for $K_{4}$ which has 7 cycles. This also shows that counts 4 and 5 are impossible.
@@
Adding one or two extra ears to a $\varTheta_{4}^{\prime}$ in a way that creates a $\widehat{K}_{4}$ subgraph only makes graphs we have already constructed from $\widehat{K}_{4}$, so we only need to consider ears that don’t create a $\widehat{K}_{4}$ subgraph.
First we add ear three. The $\varTheta_{4}^{\prime}$ is built on a $\varTheta$, whose constituent paths we call $P,Q,R$, by placing ear two on one of the paths, say $P$. Let $P=P_{ux}P_{xy}P_{yv}$, where $P_{ux}$ and $P_{yv}$ may have length 0. (See Figure 1.) Each further ear must also be on one of $P$ as in Figure 4(a2, b) or $Q$ as in Figure 4(a1), or on ear two, which is equivalent to being on $P$ nested inside ear two as in Figure 4(a2). In Figure 4(a) we get $\binom{4}{2}=10$ cycles because its cycle count is the same as that of $\varTheta_{5}$ by Lemma 6(c). In Figure 4(b) we have a necklace of three cycles (it is the upper middle graph in Figure 2); this graph has $11$ cycles. We note that adding an ear on a previously added ear does not give any different graphs from these.
@@
Now we add ear four. The four essentially different ways to do so are shown in Figure 5. Graphs (b) show all ways to have ears two, three, and four on the same constituent path of the theta graph; graphs (a) show the other ways. The four types are represented by (a1–a3) and (b1). By Lemma 6(c) graphs (a1) and (a3) have the same cycle count as $\varTheta_{6}$, that is, $\binom{6}{2}=15$. Graph (b1) is a necklace of four 2-cycles, with a total of 20 cycles. Graph (a2) has 17 cycles (in fact by Lemma 6(c) it has the same cycle count as the bottom left graph in Figure 2). It is easy to verify that no way of adding ears on ears gives any other graphs.
@@ 2. Proof for numbers beyond 89
Now we return to the proof of the main theorem. It was shown in [3] that every integer other than 2, 4, 5, 7, 8, 9, 12, 13, 14, 16, 18, 19, 22, 23, 26, 27, 29, 31, 33, 35, 38, 39, 42, 43, 46, 50, 52, 54, 60, 65, 68, 72, 77 and 89 counts the number of subtrees of a tree. (Cf. [6, Sequence A184164].) By Theorem 8, every positive integer other than those (and 1) is the number of cycles in an inseparable graph.
@@ 3. Proof for the missing numbers up to 89
Figure 7. Graphs for the $2$ exceptional cycle counts from $2$ to 89.
@@ 4. Proof of Theorem 2
The construction in Section 2 produces inseparable, planar cubic graphs for cycle counts not listed in [3] ([6, Sequence A184164]), as we noted there. This leaves us with the counts listed in [3] apart from 1, 2, 4, 5, 8, 9, 13, and 16. A computer search using plantri [2] found planar cubic graphs for these cases, with up to 14 vertices. These can be found at [4]. An example which requires 14 vertices is shown in Figure 8.
@@ 5. Consequences
Every positive integer is the cycle count of some inseparable Hamiltonian planar graph, except for $2$, 44, 55, 88, 99, and $16$.
@@
6. Conjectures
We are not aware of solutions for other classes of graphs. To motivate such research, we present the results of some computations. In all cases, it is our conjecture, in the absence of proof, that there are no more missing cycle counts.
  • (a)
    3-connected simple graphs. Cycle counts 1–6, 8–12, 16–20, 27, 30, 32–34 are missing. All other counts up to 3000 do occur.
  • (b)
    2-connected simple cubic graphs. Cycle counts 1–6, 8–13, 23–25, 27, 31–37, 40, 43–45, 48, 58–59, 62, 64, 68, 75 and 115 are missing. All other counts up to 4000 do occur.
  • (c)
    3-connected simple cubic graphs. Cycle counts 1–6, 8–13, 16–25, 27, 30–45, 48–49, 58–78, 80, 82, 88, 92, 103, 108, 110–132, 134–136, 138–139, 142–143, 150, 162, 195, 203, 208, 210–212, 214–220, 222–227, 230–233, 235, 238–239, 243, 247, 251, 255, 403, 407, 411, 415, 419, 423, 427 and 459 are missing. All other counts up to 3000 do occur.
  • (d)
    2-connected planar cubic graphs. In addition to the missing counts in item (b), this classes misses cycle counts 15–21, 29, 51, 54, 56, 67, 70, 78, 86, 90, 107, 112, 123 and 131. All other cycle counts up to 4000 do occur.
  • (e)
    3-connected planar cubic graphs. In addition to the missing counts in item (d), this class misses cycle counts 15, 29, 51, 53–57, 86, 90, 93, 85, 97–102, and many larger counts including 1182. All cycle counts from 1183 to 4000 do occur.
@@ 7. Declarations
Data for inseparable planar graphs of order at most 8 having all cycle counts up to 100 are available at [4] under the heading “Graphs with given cycle counts”. Only distributed software was used for this project. Graph generators were geng for general graphs [5], plantri for planar graphs [2] and snarkhunter for cubic graphs [1]. The tool countg was used for cycle counting [5].